# OVERVIEW ARTICLE

### v1

@@ Abstract
Training generative AI models requires extensive amounts of data. A common practice is to collect such data through web scraping. Yet, much of what has been and is collected is copyright protected. Its use may be copyright infringement. In the USA, AI developers rely on “fair use” and in Europe, the prevailing view is that the exception for “Text and Data Mining” (TDM) applies. In a recent interdisciplinary tandem-study, we have argued in detail that this is actually not the case because generative AI training fundamentally differs from TDM. In this article, we share our main findings and the implications for both public and corporate research on generative models. We further discuss how the phenomenon of training data memorization leads to copyright issues independently from the “fair use” and TDM exceptions. Finally, we outline how the ISMIR could contribute to the ongoing discussion about fair practices with respect to generative AI that satisfy all stakeholders.
@@ 1.   Introduction
Generative artificial intelligence (AI) has emerged as a transformative tool across numerous creative domains, including music, visual art, and literature. These systems operate by learning patterns, structures, and styles from large datasets, enabling them to generate novel outputs that mimic or extend human creativity. Recent advances, particularly in deep learning architectures such as transformer models and diffusion-based approaches, have significantly enhanced the quality and diversity of AI-generated music. Models like Jukebox (Dhariwal et al.,, 2020) or MusicLM (Agostinelli et al.,, 2023) as well as text-to-music AI services such as provided by the start-ups Suno AI [^1] https://suno.com and Udio [^2] https://www.udio.com/ exemplify the state of the art, producing compositions that range from stylistic imitations to innovative genre-blending works.
@@
A key factor underlying this progress is the scale and richness of the training data. High-quality, diverse datasets allow models to capture the intricacies of musical styles, genres, and techniques. However, the reliance on extensive copyrighted material raises complex legal and ethical questions. For instance, in a legal statement [^3] United States District Court for the District of Massachusetts, Civil Action No. 1:24-cv-11611-FDS, Document 28, Filed 08/01/24. Archived at https://archive.org/details/gov.uscourts.mad.272063/ gov.uscourts.mad.272063.28.0.pdf as part of a copyright lawsuit by the Record Industry Association of America (RIAA) against Suno AI, the company openly stated that “Suno’s training data includes essentially all music files of reasonable quality that are accessible on the open Internet, abiding by paywalls, password protections, and the like, combined with similarly available text descriptions.” They further argued that “It is fair use under copyright law to make a copy of a protected work as part of a back-end technological process, invisible to the public, in the service of creating an ultimately non-infringing new product.” Time will tell whether this argument will hold in court. Irregardless of the outcome, this case illustrates how it has become increasingly difficult to navigate the legal space in the context of generative AI.
@@
Addressing this issue, this paper aims to create a better understanding of the interplay between generative AI’s technical capabilities and copyright law – particularly in the context of music information retrieval (MIR) and its broader implications for creators, researchers, and industry stakeholders. It is based by a study on generative AI and copyright law in Germany and the EU conducted by the authors in 2024 and published in German as an open-access book (Dornis and Stober,, 2024). This article contains a summary of the main findings with specific implications for the MIR community. Section 2 introduces legal frameworks for training data collection and usage and discusses their applicability to training generative AI models. The text extends the scope of the original study by incorporating a discussion of the fair use doctrine in the US and a brief coverage of regulations in jurisdictions other than the EU and the US. Section 3 addresses the issue of training data memorization and discusses the resulting legal implications. This part is essentially a translated and condensed version of the original German text of the findings in our extended analysis (Dornis and Stober,, 2024). Section 4 discusses the challenge of providing a sufficiently detailed documentation of training data sources – a topic that was not covered by the original study. Here, we propose a pragmatic solution that balances the needs of rightsholders with the constraints of AI developers. Finally, we draw conclusions and provide an outlook in Section 5.
@@ 2.   Legal Frameworks for Training Data Collection and Usage
Given the fast pace at which generative AI models – in the following abbreviated as GenAI – have evolved over the last decade since the emergence of Generative Adversarial Networks and Variational Autoencoders around 2014, it is no wonder that legislature still lags behind. In fact, there are no legal frameworks yet that explicitly address GenAI. [^4] The EU DSM Directive did not address the use of copyrighted materials for GenAI training. The 2024 AI Act, albeit the lawmakers were aware of the issue, is conceived as a regulation that primarily concerns product safety and, hence, AI risks in general. See Dornis, 2025b . Companies like OpenAI or Suno AI refer to copyright exceptions for text and data mining (TDM) and “fair use” to argue that what they do is legal – such as using vast quantities of mostly unlicensed data scraped from the internet as well as public and commercial libraries. In the following, we will first outline these legal framework and compare them to each other. We will then discuss whether they are actually applicable in the context of GenAI training and finally draw conclusions for MIR research.
@@ 2.1   The TDM Exceptions in the EU
In the European Union, the Directive (EU) 2019/790 on copyright in the Digital Single Market (DSM Directive) explicitly introduces a legal framework for TDM. [^5] https://eur-lex.europa.eu/eli/dir/2019/790/oj Articles 3 and 4 of this directive establish specific exceptions for TDM activities:
Article 3 provides an exception for reproductions and extractions by research organizations [^6] A “research organisation” is either non-profit entity or has a public service research mission. and cultural heritage institutions777“Cultural heritage institutions” are publicly accessible libraries, museums, archives and film or audio heritage institution. for the purposes of scientific research. Key aspects include:
  • Lawful Access: Works must be accessible (i.e. without circumventing restrictions). This includes content under subscriptions or open access licenses.
@@
  • No Opt-Out: Rightsholders cannot prevent access through contractual terms or licenses.
@@
  • Opt-Out: Rightsholders can opt out by reserving their rights in a machine-readable manner, such as through a robots.txt file. [^9] The actual wording is “in an appropriate manner, such as machine-readable means in the case of content made publicly available online”. There is no technical standard for this yet, which leaves room for a discussion of what is “appropriate.” A common approach seems to be to put this in the robots.txt file. A recent argument is that anything a human can read and understand – such as text on an imprint page of a website – is deemed machine-readable as well.
The DSM Directive’s provisions are relatively liberal for non-commercial research, allowing activities such as scraping the internet for MIDI and audio files or using subscription services like Spotify or Deezer for data collection. This even implies that using freely downloadable pirated content is covered by the exception. However, for commercial applications, restrictions are more stringent, requiring data deletion after use and accommodating rightsholder opt-outs. Yet, use of freely available content is also generally considered admissible, even if it should have been found on pirate websites. As a result, some content owners have started to negotiate specific paid-for TDM licences with for-profit miners.
@@ 2.2   The Fair Use Doctrine in the US
In the United States, the doctrine of “fair use” under Section 107 of the Copyright Act (17 U.S.C. § 107) [^10] https://uscode.house.gov/view.xhtml?edition=prelim&num=0& req=granuleid%3AUSC-prelim-title17-section107 permits the use of copyrighted material without explicit authorization for purposes such as research, education, criticism, or commentary, provided that such use meets a four-factor test:
@@
    Purpose and Character of Use: Non-commercial and transformative uses are favored.
@@
    Amount and Substantiality: Limited use of the copyrighted material is preferred.
@@
    Effect on the Market: Uses that do not harm the market value of the original work are more likely to qualify as fair use.
Fair use provides greater flexibility compared to the EU’s DSM Directive, particularly for commercial TDM activities. While the EU imposes explicit restrictions on data retention and grants rightsholders the ability to opt out (Article 4), the US framework does not require prior authorization or data deletion as long as the use can be justified as fair. Prima facie, therefore, this broad interpretation allows commercial entities to engage in TDM without negotiating licenses, creating a competitive advantage for US-based researchers and businesses.
@@
In contrast, the DSM Directive offers a more structured approach, with clear provisions for different types of TDM activities and the roles of rightsholders. By explicitly outlining permissible activities and conditions, the EU framework reduces ambiguity and provides more predictable guidelines for compliance.
@@ 2.3   Other Global Frameworks
Canada: Canada’s Copyright Act [^11] https://laws-lois.justice.gc.ca/eng/acts/C-42/ includes fair dealing provisions that may apply. However, their application to TDM or GenAI training is underexplored and lacks specific guidelines. Legal reform is currently debated. [^12] https://ised-isde.canada.ca/site/strategic-policy-sector/en/marketplace-framework-policy/consultation-paper-consultation-copyright-age-generative-artificial-intelligence#s21
@@
China: China’s Copyright Law, as amended in 2021, does not explicitly address TDM. [^13] See, e.g., Hua, (2022). However, the “Interim Measures for the Management of Generative Artificial Intelligence Services” [^14] https://www.cac.gov.cn/2023-07/13/c_1690898327029107.htm (2023) mandate that GenAI services comply with copyright regulations. Also, the 2024 draft AI Law contains further provisions on the use of copyrighted material for GenAI training.
@@
India: India’s Copyright Act of 1957151515https://copyright.gov.in/ has no specific exceptions for TDM, but rather provides for a general rule on “fair dealing”. [^16] Section 52 Indian Copyright Act. Since this rule is interpreted rather narrowly, however, it is arguable whether it could help justify TDM and GenAI training activities. Recent lawsuits, such as Asian News International vs. OpenAI, highlight the need for clearer regulations.
@@
Japan: Japan’s Copyright Act [^17] https://www.cric.or.jp/english/clj/cl2.html seems to be one of the most permissive frameworks globally. Most notably, Article 30-4 allows the use of copyrighted works for purposes that do not involve the expressive use of the work (loosely translated as “enjoyment”) which clearly includes TDM. Yet, the actual scope of justification that is granted under Japanese law is still debated and far from clear.
@@
South Korea: Proposed amendments to South Korea’s Copyright Act [^18] https://elaw.klri.re.kr/eng_service/lawView.do?hseq=42726& lang=ENG aim to permit TDM. Yet, here as well, whether the TDM exception will cover all kinds of GenAI training is disputed and far from clear. [^19] See, e.g., Korea Copyright Commission, A Guide on Generative AI and Copyright, 2023; https://www.copyright.or.kr/eng/doc/etc_pdf/ Guide_on_Generative_AI_and_Copyright.pdf
@@
United Kingdom: The UK Copyright, Designs, and Patents Act 1988202020https://www.legislation.gov.uk/ukpga/1988/48/section/29A does provide for a TDM exception, but this exception is limited to “computational analysis” on lawfully accessed works for the “sole purpose of research for a non-commercial purpose”. Generally, it is interpreted to exclude commercial TDM. Currently, the government is considering new regulation similar to the EU.
@@
Looking at this (meagre) overview highlights both divergence and convergence: Some jurisdictions – similar to the EU – attempt to offer clear guidelines for TDM exceptions, others rather rely on more open-ended and vague fair use or fair dealing provisions. Overall, agreement seems to exist that GenAI training must not be made impossible. Yet, the silver bullet still must be invented. This status of lawmaking underscores the importance of international dialogue on harmonizing policies for TDM and GenAI, especially given the global nature of AI research and data use.

### v2

@@ Abstract
Training generative AI models requires extensive amounts of data. A common practice is to collect such data through web scraping. Yet, much of what has been and is collected is copyright protected. Its use may thus be copyright infringement. In the USA, AI developers rely on “fair use” and in Europe, the prevailing view is that the exception for “Text and Data Mining” (TDM) applies. In a 2024 interdisciplinary tandem-study, we have argued in detail that this is actually not the case because generative AI training fundamentally differs from TDM. In this article, we share our main findings and the implications for both public and corporate research on generative models. We further discuss how the phenomenon of training data memorization leads to copyright issues independently from the “fair use” and TDM exceptions. Finally, we outline how the MIR community could contribute to the ongoing discussion about fair practices with respect to generative AI that satisfy all stakeholders.
@@ 1.  Introduction
Generative artificial intelligence (AI) has emerged as a transformative tool across numerous creative domains, including music, visual art, and literature. These systems operate by learning patterns, structures, and styles from large datasets, enabling them to generate novel outputs that mimic or extend human creativity. Recent advances, particularly in deep learning architectures such as transformer models and diffusion-based approaches, have significantly enhanced the quality and diversity of AI-generated music. Models like Jukebox (Dhariwal et al., 2020) or MusicLM (Agostinelli et al., 2023) as well as text-to-music AI services such as provided by the start-ups Suno AI [^1] https://suno.com. and Udio [^2] https://www.udio.com/. exemplify the state of the art, producing compositions that range from stylistic imitations to innovative genre-blending works.
@@
A key factor underlying this progress is the scale and richness of the training data. High-quality, diverse datasets allow models to capture the intricacies of musical styles, genres, and techniques. However, the reliance on extensive copyrighted material raises complex legal and ethical questions (Goetze, 2024). For instance, in a legal statement [^3] United States District Court for the District of Massachusetts, Civil Action No. 1:24-cv-11611-FDS, Document 28, Filed 08/01/24. Archived at https://archive.org/details/gov.uscourts.mad.272063/ gov.uscourts.mad.272063.28.0.pdf. as part of a copyright lawsuit by the Record Industry Association of America (RIAA) against Suno AI, the company openly stated that “Suno’s training data includes essentially all music files of reasonable quality that are accessible on the open Internet, abiding by paywalls, password protections, and the like, combined with similarly available text descriptions.” They further argued that “It is fair use under copyright law to make a copy of a protected work as part of a back-end technological process, invisible to the public, in the service of creating an ultimately non-infringing new product.” Time will tell whether this argument will hold in court. Regardless of the outcome, this case illustrates how it has become increasingly difficult to navigate the legal space in the context of generative AI.
Our analysis builds upon a rich interdisciplinary discourse. The legal and philosophical questions raised by generative AI intersect with a longer-standing critical discourse on musical property and authorship. Scherzinger (2014) critiques the historical expansion of intellectual property rights in music, arguing that increasingly broad copyright protections contribute to the “enclosure” of shared cultural resources. Because musical creativity relies heavily on borrowing, repetition, and transformation, treating musical elements as exclusive property can paradoxically inhibit the innovation and expressive freedom that copyright is meant to encourage. This tension between protection and progress is central to the current debate on AI training data. Drott (2019) applies a similar line of analysis to the music AI industry, examining the political economy of how corporate AI development extracts value from musical works used as training data and raising fundamental questions about copyright, fair compensation, and the sustainability of a musical commons. Taking a philosophical approach to the question of AI-generated output, Caldwell (2023) examines who (or what) should be considered the “author” of AI art. Drawing on the theories of Locke, Hegel, and Kant, she argues that the end user – whose creative choices guide the system and shape the resulting work – should be treated as the author. Her position challenges inter alia the current U.S. Copyright Office’s policy denying protection to works lacking sufficient human authorship. This debate over authorship has practical consequences for documentation and cataloging, as highlighted by Berkowitz (2024). His analysis of recent U.S. Copyright Office and library cataloging guidelines reveals them internally inconsistent and poorly suited to describing human–AI co-creation, and he advocates for more nuanced cataloging practices oriented toward accurately representing the creative process. Finally, Clancy (2021) provides a comprehensive systemic perspective in his doctoral thesis, modeling the broader “music ecosystem” affected by AI technologies. He argues that existing copyright frameworks are insufficient to address the economic disruptions introduced by AI-generated music and that sustainable responses will require not only legal reform but also a foundational ethical response from all stakeholders. Collectively, these works establish that the challenges posed by generative AI are not merely technical or doctrinal, but are deeply embedded in ongoing struggles over musical value, authorial identity, and the ethical governance of cultural production – a context that underscores the need for the pragmatic documentation framework we propose.
@@
With this paper, we aim to create a better understanding of the interplay between generative AI’s technical capabilities and copyright law – particularly in the context of music information retrieval (MIR) and its broader implications for creators, researchers, and industry stakeholders. It is based on a study on generative AI and copyright law in Germany and the EU conducted by the authors in 2024 and published in German as an open-access book (Dornis and Stober, 2024). This article provides a summary of the main findings with specific implications for the MIR community. Section 2 introduces legal frameworks for training data collection and usage and discusses their applicability to training generative AI models. The text extends the scope of the original study by incorporating a discussion of the fair use doctrine in the US and a brief coverage of regulations in jurisdictions other than the EU and the US. Section 3 addresses the issue of training data memorization and discusses the resulting legal implications.Although under a legal-doctrinal lens, issues of “copying” and “reproduction” have to be considered prior to exceptions and limitations like, e.g., the Text-and-Data-Mining exception (TDM exception) or the fair use defense, this paper will start with an overview and analysis of the former in Section 2. This is due to the fact that, even if the training of a generative AI model should not result in model-internal reproductions of the training data, the development and utilization of such a model necessarily always brings up the question whether training data can be collected, compiled, or employed. Hence, the TDM exception and the fair use defense are the practically most relevant issues. Notwithstanding, of course, we will also address “memorization”, a topic that has come to dominate the debate more recently. Understanding the technical basis of “memorization” and its legal implications helps to conceive of a valid and workable approach to the collection, compilation, and documentation of training data for generative AI training. Section 4 discusses the challenge of providing a sufficiently detailed documentation of training data sources – a topic that was not covered by the original study. Here, we propose a pragmatic solution that balances the needs of rightholders with the constraints of AI developers. Finally, we draw conclusions and provide an outlook in Section 5.
@@ 2.  Legal Frameworks for Training Data Collection and Usage
Given the fast pace at which generative AI models have evolved over the last decade, it is no wonder that legislation still lags behind. In fact, there are no legal frameworks yet that explicitly address generative AI (hereafter GenAI). [^4] The EU DSM Directive did not address the use of copyrighted materials for GenAI training. The 2024 AI Act, albeit the lawmakers were aware of the issue, is conceived as a regulation that primarily concerns product safety and, hence, AI risks in general. These issues are debated in Dornis (2025b). Companies like OpenAI or Suno AI refer to copyright exceptions for text and data mining (TDM) and “fair use” to argue that what they do is legal – such as using vast quantities of mostly unlicensed data scraped from the internet as well as public and commercial libraries.
Before proceeding, we clarify the scope of “generative AI” as used in this paper. In machine learning, the term “generative model” has a broad technical meaning, encompassing any model that models the joint probability distribution of data and labels (e.g., Naive Bayes classifiers). However, the current legal and public debate centers on models that generate novel content – text, images, music – that closely resembles their training data. This includes large language models (LLMs) like GPT, diffusion models for image and audio generation, and music generation systems like Jukebox and MusicLM. Our analysis focuses on this subset, which poses the most pressing copyright challenges due to their capacity to produce expressive outputs that may compete with or replicate copyrighted works.
In the following, we will first outline the relevant legal frameworks and compare them. We will then discuss whether they are actually applicable in the context of GenAI training and finally draw conclusions for MIR research.
@@ 2.1  The TDM Exceptions in the EU
In the European Union, the Directive (EU) 2019/790 on copyright in the Digital Single Market (DSM Directive) explicitly introduces a legal framework for TDM. [^5] https://eur-lex.europa.eu/eli/dir/2019/790/oj. Articles 3 and 4 of this directive establish specific exceptions for TDM activities:
Article 3 provides an exception for reproductions and extractions by research organizations [^6] A “research organisation” is either a non-profit entity or has a public service research mission. and cultural heritage institutions777“Cultural heritage institutions” are publicly accessible libraries, museums, archives and film or audio heritage institutions. for the purposes of scientific research. Key aspects include:
  • Lawful access: Works must be accessible (i.e., without circumventing restrictions). This includes content under subscriptions or open access licenses.
@@
  • No opt-out: Rightholders cannot prevent access through contractual terms or licenses.
@@
  • Opt-out: Rightholders can opt out by reserving their rights in a machine-readable manner, such as through a robots.txt file. [^9] The actual wording is “in an appropriate manner, such as machine-readable means in the case of content made publicly available online”. There is no technical standard for this yet, which leaves room for a discussion of what is “appropriate.” A common approach seems to be to put this in the robots.txt file. A recent argument is that anything a human can read and understand – such as text on an imprint page of a website – is deemed machine-readable as well. For German case law on this issue see OLG Hamburg LG Hamburg GRUR 2024, 1710 – LAION; OLG Hamburg GRUR-RS 2025, 33887 – LAION.
The DSM Directive’s provisions are relatively liberal for non-commercial research, allowing activities such as scraping the internet for MIDI and audio files or using subscription services like Spotify or Deezer for data collection. This even implies that using freely downloadable pirated content is covered by the exception. However, for commercial applications, restrictions are more stringent, requiring data deletion after use and accommodating rightholder opt-outs. As a result, some content owners have started to negotiate specific paid-for TDM licences with for-profit miners.
@@ 2.2  The Fair Use Doctrine in the US
In the United States, the doctrine of “fair use” under Section 107 of the Copyright Act (17 U.S.C. § 107) [^10] https://uscode.house.gov/view.xhtml?edition=prelim&num=0& req=granuleid%3AUSC-prelim-title17-section107. permits the use of copyrighted material without explicit authorization for purposes such as research, education, criticism, or commentary, provided that such use meets a four-factor test:
@@
    Purpose and character of the use: Non-commercial and transformative uses are favored.
@@
    Amount and substantiality of the portion used in relation to the copyrighted work as a whole: Limited use of the copyrighted material is preferred.
@@
    Effect of the use upon the potential market for or value of the copyrighted work: Uses that do not harm the market value of the original work are more likely to qualify as fair use.
Fair use provides greater flexibility compared to the EU’s DSM Directive, particularly for commercial TDM activities. While the EU imposes explicit restrictions on data retention and grants rightholders the ability to opt out (Article 4), the US framework does not require prior authorization or data deletion as long as the use can be justified as fair. Prima facie, therefore, this broad interpretation allows commercial entities to engage in TDM without negotiating licenses, creating a competitive advantage for US-based researchers and businesses.
@@
In contrast, the DSM Directive offers a more structured approach, with clear provisions for different types of TDM activities and the roles of rightholders. By explicitly outlining permissible activities and conditions, the EU framework reduces ambiguity and provides more predictable guidelines for compliance.
@@ 2.3  Other Global Frameworks
Canada: Canada’s Copyright Act [^11] https://laws-lois.justice.gc.ca/eng/acts/C-42/. includes fair dealing provisions that may apply. However, their application to TDM or GenAI training is underexplored and lacks specific guidelines. Legal reform is currently debated. [^12] https://ised-isde.canada.ca/site/strategic-policy-sector/en/marketplace-framework-policy/consultation-paper-consultation-copyright-age-generative-artificial-intelligence#s21.
@@
China: China’s Copyright Law, as amended in 2021, does not explicitly address TDM. [^13] See, e.g., Hua (2022). However, the “Interim Measures for the Management of Generative Artificial Intelligence Services” [^14] https://www.cac.gov.cn/2023-07/13/c_1690898327029107.htm. (2023) mandate that GenAI services comply with copyright regulations. Also, the 2024 draft AI Law contains further provisions on the use of copyrighted material for GenAI training.
@@
India: India’s Copyright Act of 1957151515https://copyright.gov.in/. has no specific exceptions for TDM, but rather provides for a general rule on “fair dealing”. [^16] Section 52 Indian Copyright Act. Since this rule is interpreted rather narrowly, however, it is arguable whether it could help justify TDM and GenAI training activities. Recent lawsuits, such as Asian News International vs. OpenAI, highlight the need for clearer regulations.
@@
Japan: Japan’s Copyright Act [^17] https://www.cric.or.jp/english/clj/cl2.html. seems to be one of the most permissive frameworks globally. Most notably, Article 30-4 allows the use of copyrighted works for purposes that do not involve the expressive use of the work (loosely translated as “enjoyment”) which clearly includes TDM. Yet, the actual scope of justification that is granted under Japanese law is still debated and far from clear.
@@
South Korea: Proposed amendments to South Korea’s Copyright Act [^18] https://elaw.klri.re.kr/eng_service/lawView.do?hseq=42726& lang=ENG. aim to permit TDM. Yet, here as well, whether the TDM exception will cover all kinds of GenAI training is disputed and far from clear. [^19] See, e.g., Korea Copyright Commission, A Guide on Generative AI and Copyright, 2023; https://www.copyright.or.kr/eng/doc/etc_pdf/ Guide_on_Generative_AI_and_Copyright.pdf.
@@
United Kingdom: The UK Copyright, Designs, and Patents Act 1988202020https://www.legislation.gov.uk/ukpga/1988/48/section/29A. does provide for a TDM exception, but this exception is limited to “computational analysis” on lawfully accessed works for the “sole purpose of research for a non-commercial purpose”. Generally, it is interpreted to exclude commercial TDM. Currently, the government is considering new regulation similar to the EU.
@@
Looking at this overview highlights both divergence and convergence: Some jurisdictions – similar to the EU – attempt to offer clear guidelines for TDM exceptions, others rather rely on more open-ended and vague fair use or fair dealing provisions. Overall, agreement seems to exist that GenAI training must not be made impossible. Yet, the silver bullet still must be invented. This status of lawmaking underscores the importance of international dialogue on harmonizing policies for TDM and GenAI, especially given the global nature of AI research and data use.